Grade 11Physical Chemistry

Calculate the energy emitted when electron of 1.0 g-atom of hydrogen undergo transition giving the spectral lines of lowest energy in the visible region of it’s atomic spectra.

Mallika Chandra
10 Years agoGrade 11

1 Answer

Beenu Mathew
10 Years ago
E =hc/λ
1/λ  = R[1/n1- 1/n22]
=R[1/22  -1/32]
=(109678)*(5/36)
1/λ  = 15233.055555555555555555555555556
λ =6.7*10- cm or 6.7 * 10-7m
you can calculate E from this...how???
well
E=hc/λ
E= 3.03 * 10-19J
If N electrons show this transition in 1g atom of H then
energy release= E*N
3.03*10-19 * 6*1023
=182.5KJ
PLS APPROVE
BEST OF LUCK
THANK YOU
JEHOYESJIRA